OpenStudy (anonymous):

l 3/5x-8 l - 6= -6 solve the absolute value

OpenStudy (anonymous):

|3/(5x - 8)| = 0 thus 3/(5x - 8) = 0 no solution exists

OpenStudy (anonymous):

or, if you meant |(3/5x) - 8| - 6 = -6 then (3/5x) - 8 = 0 then 3/5x = 8 thus 40x = 3 so x = 3/40